# Test-Time Adaptation with State-Space Models

### v1

@@ Abstract
Distribution shifts between training and test data are all but inevitable over the lifecycle of a deployed model and lead to performance decay. Adapting the model can hopefully mitigate this drop in performance. Yet, adaptation is challenging since it must be unsupervised: we usually do not have access to any labeled data at test time. In this paper, we propose a probabilistic state-space model that can adapt a deployed model subjected to distribution drift. Our model learns the dynamics induced by distribution shifts on the last set of hidden features. Without requiring labels, we infer time-evolving class prototypes that serve as a dynamic classification head. Moreover, our approach is lightweight, modifying only the model’s last linear layer. In experiments on real-world distribution shifts and synthetic corruptions, we demonstrate that our approach performs competitively with methods that require back-propagation and access to the model backbone. Our model especially excels in the case of small test batches—the most difficult setting.
@@ 1 Introduction
Predictive models often have an ‘expiration date.’ Real-world applications tend to exhibit distribution drift, meaning that the data points seen at test time are drawn from a distribution that is different than the training data’s. Moreover, the test distribution usually becomes more unlike the training distribution as time goes on. An example of this is with recommendation systems: trends change, new products are released, old products are discontinued, etc. Unless a model is updated, its ability to make accurate predictions will expire, requiring the model to be taken offline and re-trained. Every iteration of this model life-cycle can be expensive and time consuming. Allowing models to remain ‘fresh’ for as long as possible is thus an open and consequential problem.
In this work, we propose State-space Test-time ADaptation (STAD), a method that delays the failure of a deployed model by performing unsupervised adaptation at test time. We perform this updating by modeling the dynamics of the parameters at a neural network’s final layer, making our approach widely applicable and computationally lightweight. Specifically, we use a state-space model (SSM) to track how the weight vectors in the final layer—where each vector corresponds to a class—evolve under distribution drift. To generate predictions for the newly-acquired batch of test points, we use the SSM’s fitted cluster means as the model’s updated parameters. We focus on natural data shifts caused by a gradually changing environment rather than noise-based shift, which has been the focus of previous work [37, 38]. Our contributions are as follows,
  • In Section 3.2, we present STAD, a state-space model to learn the dynamics of how a classifier’s last-layer weights evolve under distribution shift, without access to any labels. No previous work has explicitly modeled these dynamics, which we demonstrate is crucial via an ablation study.
  • In Sections 3.2 and 3.3, we provide two implementations of STAD—one using Gaussian distributions and one using von Mises-Fisher distributions. Each represents a different assumption about the geometry of the model’s last-layer representations.
  • In Section 5, we apply STAD to real-world (or ‘natural’) shifts, showing improvements over state-of-the-art adaptation methods such as TENT [37] and CoTTA [38]. Previous work has mostly focused on synthetic noise-based shifts, and our method is competitive on this task as well.
@@ 2 Problem Setting
We focus on the traditional setting of multi-label classification, where $\mathcal{X}\subseteq{\mathbb{R}}^{D}$ denotes the input (feature) space and $\mathcal{Y}\subseteq\{1,\dots,K\}$ denotes the label space. Let ${\mathbf{x}}$ and y be random variables and $\mathbb{P}({\mathbf{x}},{\textnormal{y}})=\mathbb{P}({\mathbf{x}})\ \mathbb{P}({\textnormal{y}}|{\mathbf{x}})$ the unknown source data distribution. We assume ${\bm{x}}\in\mathcal{X}$ and $y\in\mathcal{Y}$ are realisations of ${\mathbf{x}}$ and y. The goal of classification is to find a mapping $f_{\theta}$, with parameters ${\theta}$, from the input space to the label space $f_{\theta}:\mathcal{X}\to\mathcal{Y}$. Fitting the classifier $f_{\theta}$ is usually accomplished by minimizing an appropriate loss function (e.g. log loss). Yet, our method is agnostic to how $f_{\theta}$ is trained and therefore easy to use with, for instance, a pre-trained model that has been downloaded from the web.
Distribution Drift and Unsupervised Adaptation
In predictive modeling, a classifier will almost always become ‘stale,’ as the test conditions inevitably change from those observed during training. In fact, the model’s presence in the world can cause this change: a model that predicts when someone is likely to be infected with a contagious disease will, if effective and widely adopted, reduce the prevalence of that disease. Thus, we want our models to be ‘fresh’ for as long as possible and continue to make accurate predictions, in spite of data drift. More formally, let the data at test-time $t$ be sampled from a distribution $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})=\mathbb{Q}_{t}({\mathbf{x}})\ \mathbb{Q}_{t}({\textnormal{y}}|{\mathbf{x}})$ such that $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})\neq\mathbb{P}({\mathbf{x}},{\textnormal{y}})\ \forall t>0$. Of course, we do not observe labels at test time, and hence we observe only a batch of features ${\mathbf{X}}_{t}=\{{\mathbf{x}}_{1,t},\ldots,{\mathbf{x}}_{N,t}\}$, where ${\mathbf{x}}_{n,t}\sim\mathbb{Q}_{t}({\mathbf{x}})$ (i.i.d.). Given the $t$-th batch of features ${\mathbf{X}}_{t}$, the goal is to adapt $f_{\theta}$, forming a new set of parameters $\theta_{t}$ such that $f_{\theta_{t}}$ has better predictive performance on ${\mathbf{X}}_{t}$ than $f_{\theta}$ would have. Since we can only observe features, we assume that the distribution shift must at least take the form of covariate shift: $\mathbb{Q}_{t}({\mathbf{x}})\neq\mathbb{P}({\mathbf{x}})\ \forall t>0$. $\mathbb{Q}_{t}({\textnormal{y}}|{\mathbf{x}})$ could shift as well, but this will be impossible to detect in our assumed setting. If $\mathbb{Q}_{t}({\textnormal{y}}|{\mathbf{x}})$ does shift, it must do so modestly, as the classifier will become completely invalid if the relationship between features and labels changes to a severe degree. This setting is known as continual test-time adaptation (CTTA) [38, 15, 31, 29, 35, 12].
3 Modelling the Dynamics of Distribution Shifts in Representation Space
@@
We now present our method: the core idea is that CTTA can be done by tracking how distribution shift affect the model’s representations. We employ linear state-space models (SSMs) to capture how test points evolve and drift. The SSM’s cluster representations then serve as an adaptive classification head that evolves with the non-stationarity of the distribution drift. In Section 3.2, we first introduce the general model and then in Section 3.3, we propose an implementation that leverages the von-Mises-Fisher distribution to model hyperspherical features.
3.1 Modeling Shift in Representation Space
@@
Whereas previous work has mostly attempted to adapt all levels of a neural network, we take the alternative approach of adapting only the last layer. Let the classifier $f_{\theta}$ be a neural network having $L$ total layers. We will treat the first $L-1$ as a black box, assuming they transform the original feature vector ${\mathbf{x}}$ into a new (lower dimensional) representation, which we denote as ${\mathbf{h}}$. The original classifier then maps these representations to the classes as: $\mathbb{E}[y|{\mathbf{h}}]\ =\ \text{{softmax}}_{y}\left({\mathbf{W}}_{0}{\mathbf{h}}\right)$, where $\text{{softmax}}_{y}\left(\cdot\right)$ denotes the dimension of the softmax’s output corresponding to the $y$-th label index and ${\mathbf{W}}_{0}$ are the last-layer weights. As ${\mathbf{W}}_{0}$ will only be valid for representations that are similar to the training data, we will discard these parameters when performing CTTA, learning new parameters ${\mathbf{W}}_{t}$ for the $t$-th time step. These new parameters will be used to generate the adapted predictions through the same link function: $\mathbb{E}[y|{\mathbf{h}}]\ =\ \text{{softmax}}_{y}\left({\mathbf{W}}_{t}{\mathbf{h}}\right)$.
In the setting of CTTA, we observe a batch of features ${\mathbf{X}}_{t}$. Passing them through the model yields corresponding representations ${\mathbf{H}}_{t}$, and this will be the ‘data’ used for the probabilistic model we will describe below. Specifically, we will model how the representations change from ${\mathbf{H}}_{t}$ to ${\mathbf{H}}_{t+1}$. Operating on this last set of hidden representations has several benefits. Firstly, test-time adaptation is fast as ${\mathbf{H}}_{t}$ is relatively low dimensional and no backpropagation is necessary. Secondly, ${\mathbf{H}}_{t}$ could be provided by a foundation model accessed through a black-box API. Lastly, recall that we do not have the labels necessary to fit a complex classifier on ${\mathbf{H}}_{t}$ directly. In turn, for ${\mathbf{H}}_{t}$ to be useful, these representations must exhibit a clear structure (e.g. clustering) that reflects the classes. This assumption can lead to diagnostics that determine if we should perform CTTA or not, as we will see in the experiments.
(a) Illustration of STAD
@@
(b) Graphical Model
Figure 1: STAD adapts to distribution shifts by inferring dynamic class prototypes ${\mathbf{w}}_{t,k}$ at each test time point $t$ observing only neural representations ${\mathbf{h}}_{t,n}$. Left: STAD infers the evolution of class clusters over time on the representation space of the penultimate layer. Right: Representations are modeled with a dynamic mixture model. Latent class prototypes ${\mathbf{w}}_{t,k}$ evolve at each time step, cluster assignments ${\mathbf{c}}_{t,n}$ determine class membership of each neural representations ${\mathbf{h}}_{t,n}$.
3.2 A Probabilistic Model of Shift Dynamics
We now describe our general method for adaptive last-layer parameters. We assume that, while the representations ${\mathbf{H}}_{t}$ are changing over time, they are still maintaining some class structure in the form of clusters. Our model will seek to track this structure as it evolves over time. Using latent variables ${\mathbf{w}}_{t,k}$, we will assume each representation is drawn conditioned on $K$ latent vectors: ${\bm{h}}_{t,n}\sim p\left({\mathbf{h}}_{t}|{\mathbf{w}}_{t,1},\ldots,{\mathbf{w}}_{t,K}\right)$, where $K$ is equal to the number of classes in the prediction task. After fitting the unsupervised model, the $K$ latent vectors will be stacked to create ${\mathbf{W}}_{t}$, the last-layer weights of the adapted predictive model (as introduced in 3.1). For the intuition of the approach, see Figure 1(a). The blue red, and green clusters represent classes of a classification problem. As the distribution shifts from time step $t=1$ to $t=3$, the class clusters shift in representation space. Our SSM aims to track this movement so that classification performance can still be preserved. We now move on to a technical description.
@@ 3.2 A Probabilistic Model of Shift Dynamics
Let ${\mathbf{H}}_{t}=({\mathbf{h}}_{t,1},\dots,{\mathbf{h}}_{t,N_{t}})\in{\mathbb{R}}^{D\times N_{t}}$ denote the neural representations for $N_{t}$ data points at test time $t$. Let ${\mathbf{W}}_{t}=({\mathbf{w}}_{t,1},\dots,{\mathbf{w}}_{t,K})\in{\mathbb{R}}^{D\times K}$ denote the $K$ weight vectors at test time $t$. As discussed above, the weight vector ${\mathbf{w}}_{t,k}$ can be thought of as a latent prototype for class $k$ at time $t$. We denote with ${\mathbf{C}}_{t}=({\mathbf{c}}_{t,1},\dots,{\mathbf{c}}_{t,N_{t}})\in\{0,1\}^{K\times{N_{t}}}$ the $N_{t}$ one-hot encoded latent class assignment vectors ${\mathbf{c}}_{t,n}\in\{0,1\}^{K}$ at time $t$. The $k$-th position of ${\mathbf{c}}_{t,n}$ is denoted with ${\textnormal{c}}_{t,n,k}$ and is $1$ if ${\mathbf{h}}_{t,n}$ belongs to class $k$ and 00 otherwise. Like in standard (static) mixture models [5], the prior of the latent class assignments $p({\mathbf{c}}_{t,n})$ is a categorical distribution, $p({\mathbf{c}}_{t,n})=\text{{Cat}}(\bm{\pi}_{t})$ with $\bm{\pi}_{t}=(\pi_{t,1},\dots,\pi_{t,K})\in[0,1]^{K}$ and $\sum_{k=1}^{K}\pi_{t,k}=1$. The mixing coefficient $\pi_{t,k}$ gives the a priori probability of class $k$ at time $t$ and can be interpreted as the class proportions.
@@
We model the evolution of the $K$ prototypes ${\mathbf{W}}_{t}=({\mathbf{w}}_{t,1},\dots,{\mathbf{w}}_{t,K})$ with $K$ independent Markov processes. The resulting transition model is then:
@@
We use the notation ${\mathbf{H}}_{1:T}=\{{\mathbf{H}}_{t}\}_{t=1}^{T}$ to denote the representation vectors ${\mathbf{H}}_{t}$ for all time steps $T$ and analogously for ${\mathbf{W}}_{1:T}$ and ${\mathbf{C}}_{1:T}$. The model’s plate diagram is depicted in Figure 1(b). The representation ${\mathbf{H}}_{1:T}$ are the observed variables and are determined by the latent class prototypes ${\mathbf{W}}_{1:T}$ and the latent class assignments ${\mathbf{C}}_{1:T}$.
@@
where $y_{t,n}$ denotes a prediction sampled for the representation vector ${\mathbf{h}}_{t,n}$. Note that adaptation can be performed in an online fashion by optimizing Equation 5 incrementally considering points up to point $t$. To omit computing the complete-data log likelihood for an increasing sequence as time goes on, we employ a sliding window approach.
@@
The simplest parametric form for the transition and emissions models is Gaussian. In this case, the transition noise follows a multivariate Gaussian distribution with zero mean and global covariance $\bm{\Sigma}^{\mathrm{trans}}\in{\mathbb{R}}^{D\times D}$. The resulting model can be seen as a mixture of $K$ Kalman filters (KFs). For posterior inference, thanks to the linearity and Gaussian assumptions, the posterior expectation $\mathbb{E}_{p({\mathbf{W}},{\mathbf{C}}|{\mathbf{H}})}[\cdot]$ in Equation 5 can be computed analytically using the well known KF predict, update and smoothing equations [6, 5]. However, the closed-form computations come at a cost as they involve matrix inversions of dimensionality $D\times D$. Moreover, the parameter size scales $K\times D^{2}$, risking overfitting and consuming substantial memory. These are limitations of the Gaussian formulation making it costly for high-dimensional feature spaces and impractical in low resource environments requiring instant predictions. In the next section, we discuss a model for spherical features that elegantly circumvents the limitations of a fully parameterized Gaussian formulation.

### v2

# Temporal Test-Time Adaptation with State-Space Models

@@ Abstract
Distribution shifts between training and test data are inevitable over the lifecycle of a deployed model, leading to performance decay. Adapting a model on test samples can help mitigate this drop in performance. However, most test-time adaptation methods have focused on synthetic corruption shifts, leaving a variety of distribution shifts underexplored. In this paper, we focus on distribution shifts that evolve gradually over time, which are common in the wild but challenging for existing methods, as we show. To address this, we propose STAD, a probabilistic state-space model that adapts a deployed model to temporal distribution shifts by learning the time-varying dynamics in the last set of hidden features. Without requiring labels, our model infers time-evolving class prototypes that act as a dynamic classification head. Through experiments on real-world temporal distribution shifts, we show that our method excels in handling small batch sizes and label shift.
@@ 1 Introduction
Predictive models often have an ‘expiration date.’ Real-world applications tend to exhibit distribution shift, meaning that the data points seen at test time are drawn from a distribution that is different than the training data’s. Moreover, the test distribution usually becomes more unlike the training distribution as time goes on. An example of this is with recommendation systems: trends change, new products are released, old products are discontinued, etc. Unless a model is updated, its ability to make accurate predictions will expire, requiring the model to be taken offline and re-trained. Every iteration of this model life-cycle can be expensive and time consuming. Allowing models to remain ‘fresh’ for as long as possible is thus an open and consequential problem.
Test-time adaptation (TTA) (Liang et al., 2024; Yu et al., 2023) has emerged as a powerful paradigm to preserve model performance under a shifting test distribution. TTA performs online adaptation of a model’s parameters using only test-time batches of features. By requiring neither access to labels nor source data, TTA algorithms can be employed in resource-constrained environments, whereas related approaches such as domain generalization, domain adaptation and test-time training cannot. Most TTA methods operate by minimizing an entropy objective (Wang et al., 2021) or updating normalization parameters (Schneider et al., 2020; Nado et al., 2020; Niu et al., 2023).
Synthetically corrupted images (e.g. CIFAR-10-C) are by far the most commonly used benchmark for assessing progress on TTA—with previous work noting a lack of benchmark diversity (Zhao et al., 2023b). These shifts increase the degree of information loss over time, and well-performing TTA methods must learn to preserve a static underlying signal. In this work we focus on a distribution shift of quite a different nature: Temporal distribution shifts are due to factors endogenous to the environment and often encode structural change, not just information loss. For example, we will experiment with the functional map of the world (FMoW) benchmark, which has the goal of classifying how land is used as it is developed over time (e.g. rural to urban). Gradual structural change over time has been studied in related subfields, such as temporal domain generalization (Bai et al., 2023), yet received little attention in TTA. We show that this setting, temporal test-time adaptation (TempTTA), poses significant challenges for existing TTA methods.
To address this gap, we propose State-space Test-time Adaptation (STAD), a method that builds on the power of probabilistic state-space models (SSMs) to represent non-stationary data distributions over time. STAD dynamically adapts a model’s final layer to accommodate an evolving test distribution. Specifically, we employ a SSM to track the evolution of the weight vectors in the final layer, where each vector represents a class, as distribution shift occurs. For generating predictions on newly acquired test batches, we use the SSM’s updated cluster means as the new parameters. STAD leverages Bayesian updating and does not rely upon normalization mechanisms. As a consequence, STAD excels in scenarios where many TTA methods collapse (Niu et al., 2023), such as adapting with very few samples and handling online class imbalance. Our contributions are the following:
  • In Sec. 2, we detail the setting of temporal test-time adaptation (TempTTA), which aims to cope with shifts that gradually evolve due to variation in the application domain. Despite being ubiquitous in real-world scenarios, these shifts are understudied in the TTA literature and pose significant challenges to established methods, as we demonstrate in Sec. 5.1 (Tab. 2).
  • In Sec. 3, we propose STAD, a novel method for TempTTA. It adapts to temporal distribution shifts by modeling its dynamics in representation space. No previous work has explicitly modeled these dynamics, which we demonstrate is crucial via an ablation study (Tab. 5).
  • In Sec. 5, we conduct a comprehensive evaluation of STAD and prominent TTA baselines under authentic temporal shifts. Our results show that STAD provides persistent performance gains, particularly in the cases of label shift and small batch sizes.
@@ 2 Problem Setting
We focus on the traditional setting of multi-class classification, where $\mathcal{X}\subseteq{\mathbb{R}}^{D}$ denotes the input (feature) space and $\mathcal{Y}\subseteq\{1,\dots,K\}$ denotes the label space. Let ${\mathbf{x}}$ and y be random variables and $\mathbb{P}({\mathbf{x}},{\textnormal{y}})=\mathbb{P}({\mathbf{x}})\ \mathbb{P}({\textnormal{y}}|{\mathbf{x}})$ the unknown source data distribution. We assume ${\bm{x}}\in\mathcal{X}$ and $y\in\mathcal{Y}$ are realisations of ${\mathbf{x}}$ and y. The goal of classification is to find a mapping $f_{\theta}$, with parameters ${\theta}$, from the input space to the label space $f_{\theta}:\mathcal{X}\to\mathcal{Y}$. Fitting the classifier $f_{\theta}$ is usually accomplished by minimizing an appropriate loss function (e.g. log loss). Yet, our method is agnostic to how $f_{\theta}$ is trained and therefore easy to use with, for instance, a pre-trained model from the web.
Table 1: TempTTA compared
source data free evolving test dist. time index bounded shift
TDG
TTA
CTTA
TempTTA
Temporal Test-Time Adaptation (TempTTA)
We are interested in adapting a model at test-time to a test distribution that evolves with time. Such temporal distribution shifts have been the focus of work in temporal domain generalization (TDG) (Nasery et al., 2021; Qin et al., 2022; Bai et al., 2023; Zeng et al., 2024). Our setting differs from TDG in that we do not require access to the source data and do not alternate the training procedure. While previous work on continual test-time adaptation (CTTA) study adaptation to changing target domains, they mostly discrete domains (e.g. different corruption types). e consider a special case of CTTA that is defined by two key aspects: the domain index is temporal and shifts occur gradually over time. See Tab. 1 for a comparison. More formally, let $\mathcal{T}=\{1,\dots,T\}$ be a set of $T$ time indices. At test time, let the data at time $t\in\mathcal{T}$ be sampled from a distribution $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})=\mathbb{Q}_{t}({\mathbf{x}})\ \mathbb{Q}_{t}({\textnormal{y}}|{\mathbf{x}})$. The test distributions differ from the source distribution, $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})\neq\mathbb{P}({\mathbf{x}},{\textnormal{y}})\ \forall t>0$, and are non-stationary, meaning $\mathbb{Q}_{t}({\mathbf{x}},{\textnormal{y}})\neq\mathbb{Q}_{t^{\prime}}({\mathbf{x}},{\textnormal{y}})$ for $t\neq t^{\prime}$. However, the test distribution evolves gradually such that the discrepancy between consecutive distributions is bounded, $0\leq d(\mathbb{Q}_{t},\mathbb{Q}_{t+1})\leq\epsilon$, where $d$ is a divergence function (Qin et al., 2022). Like in standard TTA, we of course do not observe labels at test time, and hence we observe only a batch of features ${\mathbf{X}}_{t}=\{{\mathbf{x}}_{1,t},\ldots,{\mathbf{x}}_{N,t}\}$, where ${\mathbf{x}}_{n,t}\sim\mathbb{Q}_{t}({\mathbf{x}})$ (i.i.d.). Given the $t$-th batch of features ${\mathbf{X}}_{t}$, the goal is to adapt $f_{\theta}$, forming a new set of parameters $\theta_{t}$ such that $f_{\theta_{t}}$ has better predictive performance on ${\mathbf{X}}_{t}$ than $f_{\theta}$ would have. Since we can only observe features, we assume that the distribution shift must at least take the form of covariate shift: $\mathbb{Q}_{t}({\mathbf{x}})\neq\mathbb{P}({\mathbf{x}})\ \forall t>0$. In addition, a label shift may occur, which poses an additional challenge: $\mathbb{Q}_{t}({\mathbf{y}})\neq\mathbb{P}({\mathbf{y}})\ \forall t>0$.
3 Tracking the Dynamics of Temporal Distribution Shifts
@@
We now present our method: the core idea is that adaptation to temporal distribution shifts can be done by tracking its gradual change in the model’s representations. We employ linear state-space models (SSMs) to capture how test points evolve and drift. The SSM’s cluster representations then serve as an adaptive classification head that evolves with the non-stationarity of the distribution shift. Fig. 1 illustrates our method. In Sec. 3.2, we first introduce the general model and then, in Sec. 3.3, we propose an efficient implementation that leverages the von-Mises-Fisher distribution to model hyperspherical features.
3.1 Adaptation in Representation Space
@@
Following previous work (Iwasawa & Matsuo, 2021; Boudiaf et al., 2022), we adapt only the last layer of the source model. This lightweight approach is reasonable for Temp-TTA since the distribution shifts only gradually, hence constrained adaptation is needed. From a practical perspective, this circumvents backpropagation through potentially large networks such as foundation models and allows adaptation when only embeddings are provided e.g. by an API. More formally, let the classifier $f_{\theta}$ be a neural network with $L$ total layers. We will treat the first $L-1$ layers, denoted as $f_{\theta}^{L-1}$, as a black box that transforms the original feature vector ${\mathbf{x}}$ into a new (lower-dimensional) representation, which we denote as ${\mathbf{h}}$. The original classifier then maps these representations to the classes as: $\mathbb{E}[y|{\mathbf{h}}]\ =\ \text{{softmax}}_{y}\left({\mathbf{W}}_{0}{\mathbf{h}}\right)$, where $\text{{softmax}}_{y}\left(\cdot\right)$ denotes the dimension of the softmax’s output corresponding to the $y$-th label index and ${\mathbf{W}}_{0}$ are the last-layer weights. As ${\mathbf{W}}_{0}$ will only be valid for representations that are similar to the training data, we will discard these parameters when performing TempTTA, learning new parameters ${\mathbf{W}}_{t}$ for the $t$-th time step. These new parameters will be used to generate the adapted predictions through the same link function: $\mathbb{E}[y|{\mathbf{h}}]\ =\ \text{{softmax}}_{y}\left({\mathbf{W}}_{t}{\mathbf{h}}\right)$. In the setting of TempTTA, we observe a batch of features ${\mathbf{X}}_{t}$. Passing them through the model yields corresponding representations ${\mathbf{H}}_{t}$, and this will be the ‘data’ used for the probabilistic model we will describe below. Specifically, we will model how the representations change from ${\mathbf{H}}_{t}$ to ${\mathbf{H}}_{t+1}$ next.
3.2 A Probabilistic Model of Shift Dynamics
Figure 1: STAD illustrated: It adapts to distribution shifts by inferring dynamic class prototypes ${\mathbf{w}}_{t,k}$ for each class $k$ (different colors) at each test time point. It operates on the representation space of the penultimate layer.
We now describe our general method for a time-evolving adaptive classification head. We assume that, while the representations ${\mathbf{H}}_{t}$ are changing gradually over time, they are still maintaining some class structure in the form of clusters. Our model will seek to track this structure as it evolves. For the intuition of the approach, see Fig. 1. The blue red, and green clusters represent classes of a classification problem. As the distribution shifts from time step $t=1$ to $t=3$, the class clusters shift in representation space. Using latent variables ${\mathbf{w}}_{t,k}$ for the cluster centers, we will assume each representation is drawn conditioned on $K$ latent vectors: ${\bm{h}}_{t,n}\sim p\left({\mathbf{h}}_{t}|{\mathbf{w}}_{t,1},\ldots,{\mathbf{w}}_{t,K}\right)$, where $K$ is equal to the number of classes in the prediction task. After fitting the unsupervised model, the $K$ latent vectors will be stacked to create ${\mathbf{W}}_{t}$, the last-layer weights of the adapted predictive model (as introduced in Sec. 3.1). We now move on to a technical description.
@@ 3.2 A Probabilistic Model of Shift Dynamics
Figure 2: Graphical Model: Representations ${\mathbf{h}}_{t,n}$ are modeled with a dynamic mixture model. Latent class prototypes ${\mathbf{w}}_{t,k}$ evolve at each time step, cluster assignments ${\mathbf{c}}_{t,n}$ determine class membership.
@@
Let ${\mathbf{H}}_{t}=({\mathbf{h}}_{t,1},\dots,{\mathbf{h}}_{t,N_{t}})\in{\mathbb{R}}^{D\times N_{t}}$ denote the neural representations for $N_{t}$ data points at test time $t$. Let ${\mathbf{W}}_{t}=({\mathbf{w}}_{t,1},\dots,{\mathbf{w}}_{t,K})\in{\mathbb{R}}^{D\times K}$ denote the $K$ weight vectors at test time $t$. As discussed above, the weight vector ${\mathbf{w}}_{t,k}$ can be thought of as a latent prototype for class $k$ at time $t$. We denote with ${\mathbf{C}}_{t}=({\mathbf{c}}_{t,1},\dots,{\mathbf{c}}_{t,N_{t}})\in\{0,1\}^{K\times{N_{t}}}$ the $N_{t}$ one-hot encoded latent class assignment vectors ${\mathbf{c}}_{t,n}\in\{0,1\}^{K}$ at time $t$. The $k$-th position of ${\mathbf{c}}_{t,n}$ is denoted with ${\textnormal{c}}_{t,n,k}$ and is $1$ if ${\mathbf{h}}_{t,n}$ belongs to class $k$ and 00 otherwise. Like in standard (static) mixture models, the prior of the latent class assignments $p({\mathbf{c}}_{t,n})$ is a categorical distribution, $p({\mathbf{c}}_{t,n})=\text{{Cat}}(\bm{\pi}_{t})$ with $\bm{\pi}_{t}=(\pi_{t,1},\dots,\pi_{t,K})\in[0,1]^{K}$ and $\sum_{k=1}^{K}\pi_{t,k}=1$. The mixing coefficient $\pi_{t,k}$ gives the a priori probability of class $k$ at time $t$ and can be interpreted as the class proportions. Next, we formally describe how we model the temporal drift of class prototypes.
@@
Fig. 2 depicts the plate diagram of our method. We model the evolution of the $K$ prototypes ${\mathbf{W}}_{t}=({\mathbf{w}}_{t,1},\dots,{\mathbf{w}}_{t,K})$ with $K$ independent Markov processes. The resulting transition model is then:
@@
We use the notation ${\mathbf{H}}_{1:T}=\{{\mathbf{H}}_{t}\}_{t=1}^{T}$ to denote the representation vectors ${\mathbf{H}}_{t}$ for all time steps $T$ and analogously for ${\mathbf{W}}_{1:T}$ and ${\mathbf{C}}_{1:T}$. We next outline how we infer the latent class prototypes ${\mathbf{W}}_{1:T}$.
@@
where $y_{t,n}$ denotes a prediction sampled for the representation vector ${\mathbf{h}}_{t,n}$. Note that adaptation can be performed online by optimizing Eqn. 5 incrementally, considering only up to point $t$. To omit computing the complete-data log likelihood for an increasing sequence as time goes on, we employ a sliding window approach.
@@
The simplest parametric form for the transition and emissions models is Gaussian. The transition noise follows a multivariate Gaussian distribution with zero mean and global covariance $\bm{\Sigma}^{\mathrm{trans}}\in{\mathbb{R}}^{D\times D}$. The resulting model can be seen as a mixture of $K$ Kalman filters (KFs). For posterior inference, thanks to the linearity and Gaussian assumptions, the posterior expectation $\mathbb{E}_{p({\mathbf{W}},{\mathbf{C}}|{\mathbf{H}})}[\cdot]$ in Eqn. 5 can be computed analytically using the well known KF predict, update and smoothing equations (Calabrese & Paninski, 2011; Bishop & Nasrabadi, 2006). However, the closed-form computations come at a cost as they involve inverting matrices of dimensionality $D\times D$. Moreover, the parameter size scales as $K\times D^{2}$, risking overfitting and consuming substantial memory. These are limitations of the Gaussian formulation making it costly for high-dimensional feature spaces and impractical in low resource environments requiring instantaneous predictions. In the next section, we discuss a model for spherical features that circumvents these limitations.